How much carbon dioxide is there in a 3.4L balloon at SATP

1
Expert's answer
2021-11-05T00:55:49-0400

Standard Ambient Temperature and Pressure (SATP):

"T = 298.15 \;K \\

p = 1\; bar =0.986923 \;atm"

Ideal Gas Law

pV=nRT

R = 0.08206 L×atm/mol×K

"n= \frac{pV}{RT} \\

n(CO_2) = \frac{0.986923 \times 3.4}{0.08206 \times 298.15}= 0.137 \;mol"

Answer: 0.137 mol